# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v2

@@ II The C*-algebra
The equivalence between the trivial and non-trivial metric space representation of quantum mechanics becomes clear if we look at it from an algebraic point of view. In this section, we will see how the same quantum system can have both trivial and non-trivial metric space descriptions (representations), and that these descriptions are unitarily equivalent. This viewpoint helps us in the subsequent subsection to define a consistent definition of a subsystem that is independent of the tensor product decomposition of the Hilbert space.
We start with a C*-algebra of operators $\mathcal{A}$ and a positive linear functional called a state $\omega$ (with norm 1) defined on this algebra. The algebra and the state $(\mathcal{A},\omega)$ together represent the statistical data of the given quantum experiment, e.g. $A\in\mathcal{A}$ can represent the measurements that can be performed on a system, and the number $\omega(A)$ will then represent the result of such operation. A faithful representation of the algebra is an *-isomorphism to the set of bounded operators on a Hilbert space $\pi:\mathcal{A}\rightarrow\pi(\mathcal{A})\subset\mathcal{B}(\mathcal{H})$. This is possible since the set of bounded operators on a complex Hilbert space is also a C*-algebra and the star operation is naturally defined via $\pi(A)^{*}=\pi(A^{*})$. The GNS construction  [27], is the construction of the Hilbert space description of this system, which involves first constructing a $C^{*}$ algebra on which the state is faithful, i.e., has a trivial kernel. This algebra $\mathcal{U}_{\omega}$ is constructed by first identifying the kernel of the state $K_{\omega}=\{X\in\mathcal{A}\,\text{ s.t. }\,\omega(X^{*}X)=0\}$, and “quotienting out” the kernel from the algebra: $\mathcal{U}_{\omega}:=\mathcal{A}/K_{\omega}$. In other words, the elements of $\mathcal{U}_{\omega}$ consist of the equivalence classes $\psi_{A}=\{A+X\,\text{ s.t. }\,X\in K_{\omega}\}$. One can see that the functional $\omega$ is faithful on $\mathcal{U}_{\omega}$. Moreover, $\mathcal{U}_{\omega}$ is a complex vector space, a property one can check by using the fact that $AX\in K_{\omega}\,\forall\,X\in K_{\omega}$, i.e. $K_{\omega}$ is a left ideal of $\mathcal{A}$. In the next step of the GNS construction, the state functional $\omega$ is used to define an inner product on the vector space $\mathcal{U}_{\omega}$ via $\langle\psi_{A}|\psi_{B}\rangle=\omega(A^{*}B)$. With this added structure, $\mathcal{U}_{\omega}$ becomes a pre-Hilbert space which can be completed to form a Hilbert space $\mathcal{H}_{\omega}=\overline{\mathcal{U}}_{\omega}$. The bar over the algebra denotes the completion with respect to the norm induced by the inner product. Therefore, the set of vectors $\{\psi_{A}\}$ is dense in $\mathcal{H}_{\omega}$. If one defines a representation given by the following action on the states: $\pi_{\omega}(A)\psi_{B}=\psi_{AB}$, one can then derive the relation $\omega(A^{*}B)=\langle\psi_{\mathbb{1}}|\pi_{\omega}(A^{*}B)\psi_{\mathbb{1}}\rangle$, where $\mathbb{1}$ is the identity element of the algebra $\mathcal{A}$. Therefore, $\psi_{\mathbb{1}}$ is a cyclic vector of the representation $(\mathcal{H}_{\omega},\pi_{\omega})$. This construction gives us a faithful cyclic representation of the algebra $\mathcal{U}_{\omega}$, with the representation space $\mathcal{H}_{\omega}$, the map $\pi_{\omega}$, and the cyclic vector $\psi_{\mathbb{1}}$, such that $\omega(A)=\langle\psi_{\mathbb{1}}|\pi(A)\psi_{\mathbb{1}}\rangle_{\mathcal{H_{\omega}}}$.
@@
It is also a well-known fact that such a representation is unique up to a unitary equivalence. In the subsequent part of this article, we will assume that given a $C^{*}$ algebra, one can always construct a faithful cyclic representation of a system $(\mathcal{A}_{\omega},\omega)$. The state $\omega$ is associated with an operator $\rho_{\omega}$ such that $\omega(A)=tr(\rho_{\omega}\pi_{\omega}(A))$. In particular, we will demand that the operator $\rho_{\omega}$ is an element of the set $\pi_{\omega}(\mathcal{A})$, in which case it is a unique operator fixed by the faithful representation map $\pi_{\omega}$.
@@ II.1 Cyclic representations in different metric spaces
Consider two faithful cyclic representations of the system $(\mathcal{A},\omega)$: (¯$(\bar{\mathcal{H}}_{\omega},\bar{\pi}_{\omega},\bar{\Phi}_{\omega})$ and $(\mathcal{H}_{\omega},\pi_{\omega},\Phi_{\omega})$. Let us assume that the two Hilbert spaces are different metric spaces where $\mathcal{H}_{\omega}$ is endowed with metric $G$ and ¯$\bar{\mathcal{H}}_{\omega}$ endowed with metric $G^{\prime}$. Therefore $\mathcal{H}_{\omega}=\mathcal{H}_{G}$ and ¯$\bar{\mathcal{H}}_{\omega}=\mathcal{H}_{G^{\prime}}$. Since these are cyclic representations, they must be unitarily equivalent, hence there exists a unitary map [^1] By “unitary” we mean an invertible isometry defined everywhere on the Hilbert space. In the most general sense dual of a unitary is equal to its inverse $T_{U}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$ between the two Hilbert spaces whose action is given by $\Psi_{A}\rightarrow T_{U}\Psi_{A}$, where $T_{U}$ is a bounded and positive operator. Therefore we should have $\langle\psi_{A}|\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}=\langle T_{U}\psi_{A}|T_{U}\psi_{A}\rangle_{\mathcal{H}_{G}}=\langle\psi_{A}|T_{U}^{*}T_{U}\psi_{A}\rangle_{\mathcal{H}_{G^{\prime}}}$, or in other words $T_{U}^{*}T_{U}=1$. One can check that the dual operator $T_{U}^{*}$ is given by the relation $T_{U}^{*}=G^{\prime-1}T^{\dagger}G$. Combining this with the unitarity property, we have the relation $G^{\prime}=T_{U}^{\dagger}GT_{U}$. We had discussed this relation in the introduction of the metric formalism. Here we have seen that this is a consequence of the the unitary equivalence of the cyclic representation of the same underlying algebra. The operator $T_{U}$ intertwines the algebras $\pi_{G}(\mathcal{A})$ and $\pi_{G^{\prime}}(\mathcal{A})$ such that $\pi_{G}(A)=T_{U}\pi_{G^{\prime}}(A)T_{U}^{-1}\,\forall\,A\in\,\mathcal{A}$. Note that the Hermitisization operation $\eta:\mathcal{H}_{G}\rightarrow\mathcal{H}:\psi\rightarrow\eta\psi$ is a special case of such a unitary operation. Therefore, given a system $(\mathcal{A},\omega)$, it can have unitarily equivalent faithful cyclic representations in different metric Hilbert spaces and always has a trivial metric (Hermitian) description  (see also  [28]).
To summarize, by defining the unitary operation as an invertible-isometric map between two Hilbert spaces, possibly of different metric structures, it becomes clear that the unique equivalence class of the cyclic representation of the state contains its not-trivial metric space representations.
II.2 Partitions of a system
Let us have a system $(\mathcal{A},\omega)$ with cyclic representation $(H_{\omega},\pi_{\omega},\Omega_{\omega})$ and two sub-algebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. Consider the cyclic representations of the two subsystems $\{\mathcal{A}_{1},\omega_{1}\}$ and $\{\mathcal{A}_{2},\omega_{2}\}$ given by $\{\mathcal{H}_{\omega_{1}},\pi^{1}_{\omega_{1}},\Omega_{\omega_{1}}\}$ and $\{\mathcal{H}_{\omega_{2}},\pi^{2}_{\omega_{2}},\Omega_{\omega_{2}}\}$ respectively, where $\omega_{1}$ and $\omega_{2}$ are the restriction of $\omega$ on $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively. We define the Kernel space of these states as $K_{\omega_{1}}=\{X\in\mathcal{A}_{1}\,\text{and}\,\omega(X^{*}X)=0\}$ and similarly $K_{\omega_{2}}$, and use them to construct the Hilbert spaces $\mathcal{H}_{\omega_{1}}=\overline{\mathcal{A}_{1}/K_{\omega_{1}}}$ and $\mathcal{H}_{\omega_{2}}=\overline{\mathcal{A}_{2}/K_{\omega_{2}}}$. Define the product of these spaces $\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}}$ as a linear combination of the elements of the form $\psi_{A_{i}}^{1}\psi_{B_{j}}^{2}$[29]. Here we of course have $A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$. This forms a linear vector space with an inner product defined through the state $\omega(A_{i}B_{j})=\omega_{1}(A_{i})\omega_{2}(B_{j})$. Note that this is not a faithful state over $\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}}$. This is because the kernel of $\omega$ cannot be constructed solely from the kernels $K_{\omega_{1}}$ and $K_{\omega_{2}}$. One therefore defines a Hilbert space called the direct product Hilbert space by quotienting out this kernel: $\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}=(\overline{\mathcal{H}_{\omega_{1}}\odot\mathcal{H}_{\omega_{2}})/K_{\omega}}$. Under this equivalence, the elements of the form $\psi_{A_{i}}^{1}\psi_{B_{j}}^{2}$ are mapped to the elements of the form $\psi_{A_{i}}^{1}\otimes\psi_{B_{j}}^{2}\in\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$. Finally, $(\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}},\pi^{1}_{\omega_{1}}\otimes\pi^{2}_{\omega_{2}},\Omega_{\omega_{1}}\otimes\Omega_{\omega_{2}})$ is a cyclic representation of the algebra $\mathcal{A}_{1}\vee\mathcal{A}_{2}$ defined as the minimal algebra containing $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. Note that if $\pi_{\omega}(A_{i}B_{j})=\pi^{1}_{\omega_{1}}(A_{i})\otimes\pi^{2}_{\omega_{2}}(B_{j})$, one can map the elements of $\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$ to the elements of $\mathcal{H}_{\omega}$. However, if all the elements of the larger algebra $\mathcal{A}$ can be written as a linear combination of terms like $A_{i}B_{j}$, this mapping becomes an isomorphism between $\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$ and $\mathcal{H}_{\omega}$. This is possible only if the elements of the two sub-algebras commute with each other and the linear combination of the product of their elements is equal to the total algebra. This condition, purely at the level of algebras, is a necessary condition for the factorization of the Hilbert space $\mathcal{H}_{\omega}\cong\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$, which we regard as the conventional partitioning of a quantum system. However, the algebraic statement provides a more operational meaning to the partitioning of a system and is useful in situations where Hilbert space does not factorize for a multipartite system. Therefore, motivated by earlier works  [17], we provide the following definition for a bi-partition of a quantum system (generalization to a multi-partition is straightforward).
@@ II.2 Partitions of a system
    The sub-algebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ commute: $\forall\,A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$, we have $[A_{i},B_{j}]=0$.
@@
Continuing the above discussion, let us suppose that a particular cyclic representation of $(\mathcal{A},\omega)$, given by $(\mathcal{H}_{\omega},\pi_{\omega},\Omega_{\omega})$ allows a genuine bi-partition of the system into $(\mathcal{A}_{1},\omega)$ and $(\mathcal{A}_{2},\omega)$ and has the metric operator $G$, such that $\mathcal{H}_{\omega}=\mathcal{H}_{G}=\mathcal{H}_{\omega}^{1}\otimes\mathcal{H}_{\omega}^{2}$. One therefore concludes that $G$ has a tensor product structure $G=G_{1}\otimes G_{2}$ where $G_{1}$ and $G_{2}$ are the metric operators for Hilbert spaces $\mathcal{H}_{\omega}^{1}$ and $\mathcal{H}_{\omega}^{2}$, respectively. The corresponding Hermitisization map from such a Hilbert space will maintain the tensor product structure: $\eta=\eta_{1}\otimes\eta_{2}:\mathcal{H}_{G}\rightarrow\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$. Therefore, a genuine bi-partition in any metric space is equivalent to a genuine bi-partition in the trivial metric Hilbert space. Hence, we will restrict ourselves to genuine bi-partitions in the trivial metric Hilbert space.
Let us now restrict ourselves to defining subsystems in representations $(\mathcal{H}_{G},\pi_{G})$, where $G\neq G_{1}\otimes G_{2}$, while we do have $\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$, i.e. the bi-partition in the non-trivial metric Hilbert space is not a genuine bi-partition. An observable $\pi_{G}(A)$ under this representation, is also an observable in the Hilbert space $\mathcal{H}_{G^{\prime}}$ if $\pi_{G}(A)$ commutes with the unitary map $T_{U}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$. Therefore, given an observable, there can be an infinite number of metric spaces to choose from. However, if we fix a certain set of observables $\{\pi_{G}(A_{i})\}_{i}$ to probe the system, such that the only operator that commutes with all the elements of the set is a multiple of identity, then the metric space under which this set remains a set of observables is fixed  [31]. Such a set is called an irreducible set of operators as they do not have any common invariant subspace. This discussion points to the relationship between the choice of the metric operator and the choice of observables. The question then naturally arises, given a partitioning of an algebra, to what extent is the metric operator fixed? The following theorem answers this question.

### v3

@@ II The C*-algebra
An algebra is a set that is closed under summation, scalar multiplication, and multiplication operations. A C* algebra has an additional operation defined on its elements: a * operation, which is a generalization of the conjugate-transpose operation defined on the set of matrices. We start with a C* algebra $\mathcal{A}$ of operators and a positive, linear functional called a state $\omega$ (with norm 1) defined on this algebra. The algebra and the state $(\mathcal{A},\omega)$ together represent the statistical data of the given experiment on a quantum system, e.g. $A\in\mathcal{A}$ can represent the measurements that can be performed on a system, and the number $\omega(A)$ will then represent the expectation value of such operation.
GNS representation: A faithful representation of the algebra is an *-isomorphism to the set of bounded operators on a Hilbert space
$\pi:\mathcal{A}\rightarrow\pi(\mathcal{A})$ $\subset\mathcal{B}(\mathcal{H})$ (3)
such that $\text{such that }\hskip 28.45274pt\pi(A^{*})$ $=\pi(A)^{*}.$
This is possible since the set of bounded operators on a complex Hilbert space is also a C*-algebra. The conventional Hilbert space structure of the state and operators can be constructed by the Gelfand–Naimark–Segal or GNS construction, whose details are given in the appendix B. Given a system $(\mathcal{A},\omega)$, a GNS cyclic representation gives a triplet $(\mathcal{H}_{\omega},\pi_{\omega},\Phi_{\omega})$ that represents the Hilbert space, the representation map, and a vacuum/cyclic vector on $\mathcal{H}_{\omega}$ respectively, such that $\pi_{\omega}(A)|\Phi_{\omega}\rangle=|\psi_{A}\rangle$ is a general element of $\mathcal{H}_{\omega}$ and the value of the state is given by
$\omega(A)=\langle\Phi_{\omega}|\pi_{\omega}(A)|\Phi_{\omega}\rangle_{\mathcal{H}_{\omega}}$ (4)
@@
Therefore, the choice of the representation fixes the choice of the inner product. It is also a well-known fact that such a representation is unique up to a unitary equivalence. In the subsequent part of this article, we will assume that given a $C^{*}$ algebra, one can always construct a faithful cyclic representation of a system $(\mathcal{A},\omega)$. The state $\omega$ is associated with an operator $\rho_{\omega}$ such that $\omega(A)=\operatorname{tr}(\rho_{\omega}\pi_{\omega}(A))$. In particular, we will demand that the operator $\rho_{\omega}$ is an element of the set $\pi_{\omega}(\mathcal{A})$, in which case it is a unique operator fixed by the faithful representation map $\pi_{\omega}$.
@@ II.1 Cyclic representations in different metric spaces
Before exploring the non-Euclidean metric representations of a system, we would like to elucidate the notion of a “unitary” map between Hilbert spaces with different metric operators. First, we review the notion of isometry, or the norm-preserving map, between metric spaces:
Definition 1.
Isometry: Given two Hilbert spaces $\mathcal{H}_{G}$ and $\mathcal{H}_{G^{\prime}}$, the map $\mathcal{T}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}:|\psi\rangle\rightarrow T|\psi\rangle$ is an isometry if we have $\langle\psi|\psi\rangle_{G}=\langle T\psi|T\psi\rangle_{G^{\prime}}=\langle\psi|T^{*}T\psi\rangle_{G}$.
Note that the adjoint of the operator $T$ is defined through the relation $\langle T\psi|\varphi\rangle_{G}=\langle\psi|T^{*}\varphi\rangle_{G^{\prime}}$. From here one can show that $T^{*}=G^{-1}T^{\dagger}G^{\prime}$ and the map $\mathcal{T}$ is an isometry if and only if $T^{*}T=1\Leftrightarrow T^{\dagger}G^{\prime}T=G$.
$\langle T\psi|\varphi\rangle_{G}$ $=\langle\psi|T^{*}\varphi\rangle_{G^{\prime}}$ (5)
$\Rightarrow\langle T\psi|G\varphi\rangle$ $=\langle\psi|G^{\prime}T^{*}\varphi\rangle$
$\Rightarrow\langle\psi|T^{\dagger}G\varphi\rangle$ $=\langle\psi|G^{\prime}T^{*}\varphi\rangle$
$\Rightarrow T^{*}$ $={G^{\prime}}^{-1}T^{\dagger}G.$
A unitary map is an isometry which is also invertible. If $\mathcal{T}$ is an invertible map, then $T^{-1}$ exists and $T^{*}={G^{\prime}}^{-1}T^{\dagger}G=T^{-1}$, or equivalently $G^{\prime}=T^{\dagger}GT$. Therefore we have the following definition for a unitary between metric spaces:
Definition 2.
Unitary: $\mathcal{T}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}:|\psi\rangle\rightarrow T|\psi\rangle$ is a unitary map if and only if $\mathcal{T}$ is an invertible map and $G^{\prime}=T^{\dagger}GT$.
The map $\mathcal{O}\to\mathcal{O}^{\prime}=T^{-1}\mathcal{O}T$ is a unitary map between the set of operators on $\mathcal{B}(\mathcal{H}_{G})\to\mathcal{B}\mathcal{(}\mathcal{H}_{G^{\prime}})$, and it preserves the expectation value of the observable $\mathcal{O}$ with respect to any state $\psi\in\mathcal{H}_{G^{\prime}}$ under the unitary $\mathcal{T}$ as is shown below
$\langle\psi|\mathcal{O}^{\prime}\psi\rangle_{G^{\prime}}$ $=\langle\psi|G^{\prime}\mathcal{O}^{\prime}\psi\rangle$ (6)
$=\langle\psi|G^{\prime}T^{-1}\mathcal{O}T\psi\rangle$
$=\langle\psi|T^{*}G\mathcal{O}T\psi\rangle$
$=\langle T\psi|G\mathcal{O}T\psi\rangle=\langle T\psi|\mathcal{O}T\psi\rangle_{G}.$
Finally, we note that the Hermitisization operation $\mathcal{H}_{G}\rightarrow\mathcal{H}:|\psi\rangle\rightarrow\eta|\psi\rangle$ is a special case of such a unitary operation which satisfies $\eta^{*}=G^{-1}\eta=\eta^{-1}$. Therefore, the Hermitisization map $\mathcal{O}\to\eta\mathcal{O}\eta^{-1}=\mathcal{O}_{\eta}$ is also a unitary map between the operators on $\mathcal{H}_{G}$ and $\mathcal{H}$. Note that $\eta T$ and $\eta^{\prime}$ are two different unitary maps from $\mathcal{H}_{G^{\prime}}$ to $\mathcal{H}$ (see fig. 1). Therefore, the two states $\eta T|\psi\rangle$ and $\eta^{\prime}|\psi\rangle$ are related by a unitary on $\mathcal{H}$: $\eta^{\prime}=U\eta T$. Therefore, the operators $T$ can be parametrized by the unitary $U$ on $\mathcal{H}$, and we denote them by $T_{U}$.
Figure 1: Unitary maps between different metric Hilbert spaces. The state $\psi$ in $\mathcal{H}_{G^{\prime}}$ can be mapped to a state in $\mathcal{H}$ either through Hermitisization ($\eta^{\prime}$) or through the Hilbert space $\mathcal{H}_{G}$ ($\eta T_{U}$). The resulting states are related by a unitary $U$ in $\mathcal{H}$, which implies $\eta^{\prime}=U\eta T_{U}$
Armed with this generalized notion of unitary maps, we see that a system represented by $(\mathcal{A},\omega)$ can have Euclidean as well as non-Euclidean metric space representation:
Remark: The representations $(\mathcal{H},\pi)$, $(\mathcal{H}_{G},\pi_{G})$ and $(\mathcal{H}_{G^{\prime}},\pi_{G^{\prime}})$ are all different cyclic representations of the same system if the representations are related by the following relations for all $A\in\mathcal{A}$
$\pi_{G}(A)=\eta^{-1}\pi(A)\eta=T_{U}\pi_{G^{\prime}}(A)T_{U}^{-1},\forall\ A\in\mathcal{A}.$ (7)
III Partitions of a system
We are now in a position to understand the decomposition of a system into subsystems independent of the chosen representation. The simplest example of a composite quantum system is a pair of distinguishable spin $\dfrac{1}{2}$ particles. In section IV.1 we will discuss the decomposition of this system in a metric space using the results from this section.
Let us have a system $(\mathcal{A},\omega)$ with cyclic representation $(H_{\omega},\pi_{\omega})$ and two *-subalgebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. Let $\omega_{1}$ and $\omega_{2}$ be the restriction of $\omega$ on $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively. Consider the cyclic representations of the two subsystems $\{\mathcal{A}_{1},\omega_{1}\}$ and $\{\mathcal{A}_{2},\omega_{2}\}$ given by $\{\mathcal{H}_{\omega_{1}},\pi^{1}_{\omega_{1}}\}$ and $\{\mathcal{H}_{\omega_{2}},\pi^{2}_{\omega_{2}}\}$ respectively. Now, the elements of the form $A_{i}B_{j}$, where $A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$, can be represented (appendix B for more details) through the representation $(\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}},\pi^{1}_{\omega_{1}}\otimes\pi^{2}_{\omega_{2}})$ where
$\pi^{1}_{\omega_{1}}\otimes\pi^{2}_{\omega_{2}}(A_{i}B_{j}):=\pi^{1}_{\omega_{1}}(A_{i})\otimes\pi^{2}_{\omega_{2}}(B_{j}).$ (8)
The set of operators $\{A_{i}B_{j}\}$ forms a subset of the algebra $\mathcal{A}_{1}\vee\mathcal{A}_{2}$ defined as the minimal algebra containing $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. The two sets coincide if and only if the algebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ commute: $[A_{i},B_{j}]=0$ for all $i$ and $j$. In that case $(\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}},\pi^{1}_{\omega_{1}}\otimes\pi^{2}_{\omega_{2}})$ is a faithful cyclic representation of the system $\{\mathcal{A}_{1}\vee\mathcal{A}_{2},\omega_{12}\}$ , where the state $\omega_{12}$ is defined as
$\omega_{12}(A_{i}B_{j})=\omega_{1}(A_{i})\omega_{2}(B_{j}).$ (9)
Finally, if $\mathcal{A}_{1}\vee\mathcal{A}_{2}=\mathcal{A}$, then $(\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}},\pi^{1}_{\omega_{1}}\otimes\pi^{2}_{\omega_{2}})$ is a faithful representation of $(\mathcal{A},\omega)$ where $\mathcal{H}_{\omega}\cong\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$ and $\pi_{\omega}(A_{i}B_{j})=\pi^{1}_{\omega_{1}}(A_{i})\otimes\pi^{2}_{\omega_{2}}(B_{j})$.
Therefore, at the level of algebras, the necessary condition for the factorization of the Hilbert space $\mathcal{H}_{\omega}\cong\mathcal{H}_{\omega_{1}}\otimes\mathcal{H}_{\omega_{2}}$ is that there are commuting subalgebras whose sum generates the total algebra. This provides a more operational meaning to the partitioning of a system and is useful in situations where Hilbert space does not factorize for a multipartite system. Motivated by earlier works  [17], we provide the following definition for a bi-partition of a quantum system (generalization to a multi-partition is straightforward).
@@ III Partitions of a system
    The subalgebras $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ commute: $\forall\,A_{i}\in\mathcal{A}_{1}$ and $B_{j}\in\mathcal{A}_{2}$, we have $[A_{i},B_{j}]=0$.
@@
Continuing the above discussion, let us suppose that a particular cyclic representation of $(\mathcal{A},\omega)$, given by $(\mathcal{H}_{\omega},\pi_{\omega})$ allows a genuine bi-partition of the system into $(\mathcal{A}_{1},\omega)$ and $(\mathcal{A}_{2},\omega)$ and has the metric operator $G$, such that $\mathcal{H}_{\omega}=\mathcal{H}_{G}=\mathcal{H}_{\omega}^{1}\otimes\mathcal{H}_{\omega}^{2}$. One therefore concludes that $G$ has a tensor product structure $G=G_{1}\otimes G_{2}$ where $G_{1}$ and $G_{2}$ are the metric operators for Hilbert spaces $\mathcal{H}_{\omega}^{1}$ and $\mathcal{H}_{\omega}^{2}$, respectively. The corresponding Hermitisization map from such a Hilbert space will maintain the tensor product structure: $\eta=\eta_{1}\otimes\eta_{2}:\mathcal{H}_{G}\rightarrow\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$. Therefore, a genuine bi-partition in any metric space is equivalent to a genuine bi-partition in the Euclidean metric Hilbert space. Hence, we will restrict ourselves to genuine bi-partitions in the Euclidean metric Hilbert space.
Partitioning: Let us now restrict ourselves to defining subsystems in representations $(\mathcal{H}_{G},\pi_{G})$, where $G\neq G_{1}\otimes G_{2}$, while we do have $\mathcal{H}_{\mathbb{1}}=\mathcal{H}^{1}\otimes\mathcal{H}^{2}$, i.e. the bi-partition in the non-Euclidean metric Hilbert space is not a genuine bi-partition. An observable $\pi_{G}(A)$ under this representation is also an observable in the Hilbert space $\mathcal{H}_{G^{\prime}}$ if $\pi_{G}(A)$ commutes with the unitary map $T_{U}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}$. Therefore, given an observable, there can be an infinite number of metric spaces to choose from. However, if we fix a certain set of observables $\{\pi_{G}(A_{i})\}_{i}$ to probe the system, such that the only operator that commutes with all the elements of the set is a multiple of identity, then the metric space under which this set remains a set of observables is fixed  [40]. Such a set is called an irreducible set of operators as they do not have any common invariant subspace. This discussion points to the relationship between the choice of the metric operator and the choice of observables. The question then naturally arises, given a partitioning of an algebra, to what extent is the metric operator fixed? The following theorem answers this question.